OpenStudy (anonymous):

If tanθ = 5/12, find secθ

OpenStudy (wolf1728):

Look up arc tangent of 5/12. Then look up the secant of this angle.

OpenStudy (anonymous):

Use the trig identity: 1 + tan^2 x = sec^2 x Replace tan x = 5/12 --> (1 + tan^2 x) =1 + 25/144 = 1.174 sec x = ?